OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half.

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\]with \[x_1=6,y_1=-3,m=\frac{1}{2}\] make a direct substitution

OpenStudy (anonymous):

i know i already did that @satellite73

OpenStudy (anonymous):

What i don't get is the answer choices y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (anonymous):

that is the point slope form

OpenStudy (anonymous):

if you already did it you have exactly one of your answer choices

OpenStudy (anonymous):

Ohhh its c Right?